You purchased 250 shares of General Motors stock at a price of $89.79 two years ago. You sold all stocks today for $87.95. During this period the stock paid dividends of $5.15 per share. What is your annualized holding period return (annual percentage rate)?

Round the answers to two decimal places in percentage form. (Write the percentage sign in the "units" box)

Your Answer:
Answer #1

Initial Investment = 89.79 * 250

= 22447.5

Value at the end of the period = 87.95 * 250

= 21987.5

Dividend Income = 5.15 * 250

= 1287.5

Holding Period Return = [Income + ( Value at the end - Initial Investment)] / Initial Investment

= [1287.5 + ( 21987.5 - 22447.5)] / 22447.5

= ( 1287.5 - 460) / 22447.5

= 3.686%

Holding Period return is  3.686%

Annualized Holding period return = [( 1 + Holding period return) ^ ( 1 / number of year)] - 1

=[ ( 1 + 3.686%) ^ ( 1/2)] - 1

= 1.8263%

Annualized Holding period return is 1.8263%